OpenStudy (anonymous):

Double check my work? I have to find the simplified expression and think I got the wrong answer. I will put the equation in the comments

OpenStudy (anonymous):

\[5\sqrt{v^8} -\sqrt{v^3} - \sqrt{16v^8} + 3\sqrt{v^3}\]

OpenStudy (anonymous):

\[v^8=(v^2)^4\]\[v^3=v^{2+1}\]\[brake~~~~the~~~~the~~~~roots.\]

OpenStudy (anonymous):

Any ideas?

hero (hero):

Keep in mind a couple things: \[\sqrt{ab} = \sqrt{a}\sqrt{b}\] In this case, \(\sqrt{16v^8} = \sqrt{16}\sqrt{v^8}\) Also, pair the terms in this manner: \[5\sqrt{v^8} - \sqrt{16v^8} + 3\sqrt{v^3}-\sqrt{v^3}\]

hero (hero):

Remember: \(\sqrt{16} = 4\) \(\sqrt{(x^a)^2} = x^a\)

hero (hero):

In this case: \(\sqrt{v^8} = \sqrt{(v^4)^2} = v^4\)

hero (hero):

That might seem like a lot to absorb at once. There's one last thing I want to mention.

hero (hero):

Distributive Property: \(ab + ac = a(b + c)\) You may be wondering how distributive property is relevant to this.

hero (hero):

Well...notice that \(\sqrt{v^8}\) is common to two terms and \(\sqrt{v^3}\) is common to two terms.

hero (hero):

We can factor those out using distributive property. But first we should reduce the expressions in accordance with the aforementioned rules.

hero (hero):

\(5\sqrt{v^8} - \sqrt{16v^8} + 3\sqrt{v^3}-\sqrt{v^3}\) becomes: \(5\sqrt{v^8} - 4\sqrt{v^8} + 3\sqrt{v^3}-\sqrt{v^3}\) by square root property \((5 - 4)\sqrt{v^8} + (3 - 1)\sqrt{v^3}\) by distributive property \((1)\sqrt{v^8} + (2)\sqrt{v^3}\) by simplification

hero (hero):

The only thing left to reduce is \(\sqrt{v^8}\) We already know from above that it simplifies to \(v^4\) so the final simplified resullt is: \(v^4 + 2\sqrt{v^3}\)